Calculate the dot product of C= (-2.-6)N and D= (-2,3)N.


1
Expert's answer
2022-01-11T09:04:29-0500
"C.D=(C_1\times D_1)+(C_2\times D_2),""C.D=(-2\ N\times(-2\ N))+(-6\ N\times3\ N)=4\ N^2-18\ N^2=-14\ N^2."